OpenStudy (anonymous):

If you wanted to eliminate the y variable by adding the two equations in the following system, what could be your first step? 5x − 6y = 6 x − y = 12

OpenStudy (anonymous):

You would need to take the bottom equation (x-y =12) and multiply it by -6 so that when you add them they cancel each other out.

OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 4x + y = 8 x − 2y = 2

OpenStudy (anonymous):

can you help me with that?

OpenStudy (anonymous):

i need to know what x is

OpenStudy (anonymous):

4x + y = 8 x - 2y = 2 (-4) which gives you -4x + 8y = -8 Then you will add the two equations together. x is eliminated and it becomes 9y = 0 so y = 0. Plug that in x -2(0) = 2 and solve so x = 2.

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

Yep